Ok, so here is the word problem, can you tell me if I set it up correctly?!
Jack bought a new camera but when it came time to record the purchase in his check register all he could remember was that the price was:
six (6) less than the square of this amount is equal to the amount itself.
So the equation should look like x^2-6=x
Am I correct?

Answer by Theo(13342)

you are correct.
x = x^-6
solving that, we get x = 3 as follows:
x = x^2 - 6
subtract x from both sides to get:
x^2 - x - 6 = 0
which becomes:
(x-3)*(x+2) = 0
which means that:
x = 3 or x = -2
since x can't be negative, then x = 3 is your only possible solution.
-----
3 = 9-6 = 3 is good, so x = 3 is your answer and your equation of:
x^ - 6 = x
is good.
-----